### v1

@@ 6.1.1 Data manipulation
Almahfouz Nasser et al. [5] proposed a Pre-processing Homogenizer strategy for reducing domain shift. Their pre-processing pipeline consists of training an auto-encoder to reconstruct the input images using a mean square loss. The main idea is to learn a uniform domain appearance of input images. To further erase domain specific signals, they incorporated an adversarial loss to the encoder using a domain discriminator. The domain discriminator is a network trained to detect the domain label (i.e., the scanner technology used to acquire images) using features arising from the encoder. In contrast, the autoencoder is trained to maximize the domain label prediction loss and minimize the reconstruction loss simultaneously to obtain domain-invariant features. This solution was applied for the MIDOG 2021 challenge (Table 1). Unfortunately, according to the leaderboard on grand challenge, this method performed very poorly with a F1 score of 0.0030 on the test set.
Yin et al. [224] proposed a novel DG method based on Adversarial Frequency Alignment (AFA). Their DG framework comprises four networks: an encoder-decoder in the Adaptive Transition Module (ATM), a fidelity discriminator, a multi-domain discriminator and a classification network. The goal of the ATM module (autoencoder) is to learn a frequency attention map that can align different domain images in a common frequency domain. It can be considered as a "data homogenizer". Given an input image coming from multiple source domains, it produces a reconstructed image, which eliminates significant domain shifts in the frequency domain. To this end, it is trained using adversarial learning by employing a fidelity discriminator and a multi-domain discriminator in the frequency domain. In contrast, the classification model is trained in the original spacial domain to classify nodules. It is applied after the ATM module. This method was evaluated in the context of lung nodule detection from CT images. The evaluation on both LUNA-DG and on an in-house dataset showed that the AFA is an efficient method. They reported a competition performance metric [138] of 0.911 on the target test set of LUNA-DG.
@@
Gunasinghe et al. [70] also investigated data homogenization to enhance the performances of DL models when applied to data with variations stemming from the use of different medical devices. Three different preprocessing methods were considered: median filter, input standardization, and randomized multi-image histogram matching. The median filter is a non-linear digital preprocessing technique, used to remove noise from an image. Input standardization aims to attenuate illumination variations. In their work, they used a method inspired by Quellec et al. [149], where RGB images are converted to the YCbCr color space and the channel holding the luminance information (Y) is standardized. Following this normalization, the color space is changed back to RGB. Histogram matching is a technique that transforms the histograms of the red, green and blue channels of an image to to match those of a specific reference image. In randomized multi-image histogram matching, histogram matching is performed sequentially using multiple reference images selected from the training source domain. The authors studied these preprocessing strategies in combination and individually to evaluate their impact on the model generalization for glaucoma detection using fundus photographs. To this end, the experiments were conducted on two public datasets, RIMONEv2 and REFUGE which include different fundus camera devices. The results have shown that standardization of images led to greater performances in most scenarios with an average of 0.85 Area Under the Receiver Operator Characteristic Curve (AUC).
Garrucho et al. [59] presented another data homogenization strategy, named intensity scale standardization. Intensity along with texture are considered to be domain-dependent feature. Inspired by Nyúl et al. [140], they proposed to perform intensity scale standardization for attenuating domain shift. It is a two-step technique: 1) it learns a standardized histogram from the training images, extracting the histogram landmarks, and 2) these landmarks are used to linearly map the intensities of input images. The authors applied this technique for mass detection in mammography. They showed that it enhanced the generalization performance. It outperformed MixStyle, Cutout, RandConv and histogram equalization.
@@
Wang et al. [205] introduced a method for harmonizing data from different sources. Inspired by Pawlowski et al. [144], they propose to use normalizing-flow-based method for counterfactual inference within a Structural Causal Model (SCM), to attain harmonization of data. The idea is to explicitly model the causal relationship of known confounders such as site, gender and age, and ROI features (i.e., the imaging measurement) in a SCM which uses normalizing flows to model probability distributions. Counterfactual inference can be performed upon such a model to sample harmonized data by intervening upon these variables. This method has been evaluated for age regression and Alzheimer’s disease classification. It showed better cross-domain generalization compared to state-of-the-art algorithms such as ComBat and IRM, and to models trained on raw data.
@@
In this context, Li et al. [115] proposed a Fourier-based method that generates adversarial data via spontaneous Amplitude SPECTrum diverSification (ASPECTS) for single-DG. This method is inspired by a multi-source DG [215] which improves the diversity of training data by combining the frequency components of samples from different source domains. Specifically, the suggested method first applies the Discrete Fourier transform to convert the image into the frequency domain, and then generates diverse samples by modifying the amplitude spectrum using a variety of randomization operations. The premise of the suggested method is that the amplitude spectrum of images after Fourier transformation mostly contains domain-invariant information. Therefore, the authors proposed to generate new adversarial samples by modifying the amplitude spectrum, i.e., randomize the amplitude and position of points in the amplitude spectrum using rescaling and pixel shuffling operations. In contrast to the existing single-DG methods, their proposed method does not require to train an extra network/module for adversarial sample generation. It has been validated on two medical datasets: a public Mitosis Domain Generalization (MIDOG) dataset and a private multicenter colposcopic image datasets. The authors reported an average accuracy over all out-of-domain data of 0.6285 for the MIDOG dataset and of 0.6287 for the multicenter colposcopic image dataset.
Zhang et al. [231] integrated a similar strategy, using domain randomization in their framework for enhancing DL generalizability. The end goal was to encourage DL model to learn stable representations. Formally, a domain randomization function is applied to a given input image to produce a modified image. The modified image should retain the same high-level semantics as the original image but its superficial distribution is different. This function is implemented as amplitude mix, where the superficial distributions of an image is perturbed through linearly interpolating its amplitude spectrum with that of another image. Alternatively, this function can be implemented as color jitter where variation are introduced in terms of hue, saturation and contrast distributions. The feature stability is an indicator of domain invariance, as domain randomization keeps the high-level semantics unchanged. It is defined as the sum of the channel-wise cosine distance between the original feature and its domain-randomized counterpart.
@@
Lucieri et al. [126] presented Amplitude-Focused Amplitude-Phase Recombination for pair samples (AF-APR-P) [^3] https://github.com/adriano-lucieri/shape-bias-in-dermoscopy. It was proposed as a variation of the original Amplitude-Phase Recombination for Pair samples (APR-P)[31] approach to improve the robustness and performance of DL models. In APR-P, an image is augmented by replacing its amplitude spectrum with the spectrum of another randomly selected image from the batch. The main idea is to improve the robustness of the DL model by reducing the dependence on the amplitude spectrum and enhancing the ability to capture phase spectrum. In contrast, AF-APR-P aims to enhance the model’s ability to generalize by focusing on the amplitude spectrum of the images while altering the phase spectrum. This is achieved by swapping the phase spectrum among images but retaining their original amplitude spectrum. The authors showed improved performance on binary skin lesion classification tasks on the The International Skin Imaging Collaboration (ISIC) dataset [^4] https://www.isic-archive.com/ and the seven-point checklist criteria dataset [96].
@@
Wang and Xia [203] introduced a Domain-Ensemble Learning with Cross-domain Mixup (DELCOM) for multi-source DG. Their proposed method collaboratively train multiple domain-specific models using the data from both source and virtual domains. To this end, they constructed virtual samples, which form a virtual mixed domain that does not belong to any source domain. In particular, they extended the conventional mixup to cross-domain mixup to create a virtual domain based on the data from source domains. The original mixup technique produces convex combinations of pairs of images and their labels: it interpolates pairs of samples from the same domain that are drawn at random. In cross-domain mixup, one combine pairs of samples from different domains, to form a virtual domain $\mathcal{S}_{mix}$ that comprises virtual images ($x_{mix}$) and labels ($y_{mix}$), as formulated in the following equations:
@@
Experiments performed on chest X-rays datasets for the diagnosis of thoracic diseases showed that the proposed DELCOM method outperformed ERM and six other DG approaches (CrossGrad [170], Deep Domain-Adversarial Image Generation (DDAIG) [236], Domain adaptive ensemble learning (DAEL) [237], MixStyle [237] and Exact Feature Distribution Matching (EFDMix) [232]).
@@
Garrucho et al. [59] also studied different augmentations techniques for DG. Namely, Cutout [47], RandConv [216] and MixStyle [237]. Cutout removes square patches from training images at random locations. RandConv produces images with random local textures but consistent shapes using linear filtering. MixStyle mixes the feature statistics of two instances to synthesize new domains during the mini-batch training. In addition to these augmentation strategies, they investigated a data homogenization approach, the intensity scale standardization approach (presented in Section 6.1.1). They conducted experiments for mass detection in mammography across multiple centers. Two backbones were considered: Swin Transformer and ResNet-50. They evaluated the performances of their model using one or a combination of data augmentation strategies. The results showed that the combination of intensity scale standardization and cutout data augmentation led to the best results in all unseen domains in terms of average AUC for both backbones.
@@
To enhance their model’s generalizability to different devices, Lafarge and Koelzer [103] incorporated an extensive and aggressive data augmentation technique. They applied a sequence of transformations such as transposition, color shift, Gamma correction, Hue rotation, spatial shift, additive Gaussian noise and cutout [47]. They applied this method for mitotic figure detection in WSI from MIDOG challenge. The evaluation on the preliminary test set of the challenge resulted in a F1 score of 0.6828.
@@
Dexl et al. [48] performed a single augmentation to each image as part as a very simple random augmentation approach, inspired by Trivial Augment [137]. These augmentations are uniformly selected from a pool of color, noise, and special transformations, with the intensity of augmentation randomly chosen within a predefined range. The pool of augmentations includes color jitter, Histogram Equalization (HE), fancy PCA, hue modification, saturation modification, equalization modification, random contrast, auto-contrast, contrast limited adaptive histogram equalization, solarize-add, sharpness, Gaussian blur, posterize, cutout, ISO noise, JPEG compression artifacts, pixel-wise channel shuffle and Gaussian noise. Each image is also randomly flipped, and the RGB channels are shuffled at random. Using this strategy, their model achieved an F1 score of 0.7138 on the preliminary test phase of the MIDOG challenge.
@@
In the domain of histopathology, conventional stain normalization methods translate pathology images into a single target stain color style, which may limit the robustness of the detection model when confronted to new stain color appearance. To address this problem, Long et al. [123] proposed to broaden the spectrum of stain color appearances in the training images for boosting the performances of the detection model to unseen data. This is achieved by introducing randomness in selecting stain normalization techniques and target color styles. In particular, they incorporated two stain normalization techniques: Reinhard [156] and Vahadane [194]. Each technique is applied using a specific probability. Vahadane transfers color in accordance with the desired color appearance matrix, whilst Reinhard modifies color depending on the target mean and variance. Using the entire training set, they derive an initial range for target mean, variance, and each component of the color appearance matrix, respectively. During augmentation, target values are chosen at random from these ranges, allowing to produce images with a wide variety of color schemes. To achieve robust detection performance for variety of images, they gradually expand those ranges to provide additional training samples to the network until there is a degradation in the detection performance. Their model, Domain Adaptive Cascade R-CNN (DAC R-CNN), achieved an F1 score of 0.7500 on the preliminary test phase of the MIDOG challenge.
Chung et al. [39] proposed a GAN-based approach to expand the style variance of the training data. Each device type is considered as a domain that produces images with distinct ’styles’, such as differences in hue and contrast. With the aim to develop a method to be robust on device types, they proposed to use style transfer using StarGAN. The generator of StarGAN [36], also referred to as the transfer module, is used to translate images into a variously styled image based on the mixing of the device characteristics, without losing morphological information (i.e., structure of a cell and nucleus) upon training. After training their transfer module, a detection network is trained using the translated images for the targeted task (mitotic figure detection). Their model achieved an F1 score of 0.7548 on the preliminary test phase of the MIDOG challenge.
Scalbert et al. [163] introduced a novel Test-Time data Augmentation (TTA) based on multi-domain image-to-image translation, using a more recent version of StarGAN (in contrast to Chung et al. [39]), StarGANV2 [37]. This TTA technique is based on the projection of images from unseen protocol into each source domain prior to classification, followed by an ensemble of predictions. They trained a StarGANV2 model, which a multi-domain image-to-image translation model, that can translate an input image $x$ into any source domain without the need of the domain label of the image. The generator of StarGANV2 is given as inputs an image and a domain-dependent style vector. These vectors can either be drawn from a random latent code or from a reference image. The authors proposed to use the former method due to the lighter size and lower computational overhead. In order to preserve the structure of the input image, the perceptual domain invariant loss is used to train the StarGANv2. After training StarGANv2 and a classifier on the source data, they used TTA at inference. Specifically, the generator of StarGANv2 projects the image into each source domain. Afterwards, the predictions of the classifier and the score of the StarGANv2 discriminator can be used to obtain a final prediction. Different strategies were explored for ensembling: 1) a naive ensembling where only the predictions of the classifier are average 2) a discriminator-based top-k ensembling, where a domain discriminator score indicates how well the test image aligns with the respective source domain. Accordingly, the top $k$ predictions linked to images with the top-k domain discriminator are combined 3) a discriminator-based weighted ensembling, where weighting the top $k$ predictions is proposed based on their domain discriminator scores to reduce the influence of poorly projected images. The proposed method has shown good results when evaluated for two different histopathology tasks: 1) patch classification of lymph node section WSIs and 2) tissue type classification in colorectal histological images. The results of their approach were superior to those using standard/ Hematoxylin&Eosin (H&E) specific color augmentation/normalization and standard test-time augmentation techniques. The naive ensembling technique showed slightly better performances than the two other proposed ensembling strategies.
@@
With the aim to improve the generalization capability of DL model to various vendors, Li et al. [118] also used a GAN-based approach for style transfer. They used CycleGAN to map the images of source domain to a vendor-style domain. The augmented images are then used in their contrastive learning strategies to develop a representation with better generalization capability to various vendor domains (Section 6.1.2). In addition, for enhancing the sample diversity, they used different diversifying operations including random cropping, random rotation, horizontal flipping, and adjustment of brightness, contrast, and saturation.
On the other hand, instead of exploiting image-to-image translation models at training phase, Yamashita et al. [219] proposed to use it at testing phase. In particular, they presented a Style TRansfer Augmentation for Histopathology (STRAP)  [^5] https://github.com/rikiyay/style-transfer-for-digital-pathology, a data augmentation that employs random style transfer from non-medical style source like paintings, with the aim to achieve domain-agnostic visual representations in computational pathology. Through style transfer, the low-level texture content of an image (which is hypothesized to be domain-specific) is changed to adopt the uninformative style of a randomly selected style source image, yet its high-level semantic content (which is supposed to be more domain-invariant) remain intact.
@@
In STRAP, the style of medical images (i.e., histopathology images), namely the texture, color and contrast are translated with the style of a selected non-medical image. However, the semantic content of the image, the global object shapes are unchanged. This is done by applying AdaIn style transfer [83], as in Geirhos et al. [60]. AdaIn style transfer creates an output image by recombining the content of an image and style of an another input image. The content and style images are encoded into a feature space by using an encoder, then the encoded representations are inputted to an AdaIN layer, which aligns the mean and variance of the content feature maps to those of the style feature maps and produces the target feature maps. The augmented image is then created from this target feature maps via a decoder. As style sources, they used different datasets, such as artistic painting from the Kaggle’s Painter by Numbers dataset [^6] https://www.kaggle.com/c/painter-by-numbers, natural images from the miniImageNet dataset [199], and the original Stanford-CRC dataset. They applied their method into two different problems: 1) colorectal cancer classification into two distinct genetic sub-types based on WSI in a single-DG setting and 2) identifying the presence or absence of breast cancer metastases in image patches extracted from histopathlogic scans of lymph node sections in a multi-source DG setting. Their method achieved higher performances compared to stain normalization based approaches. For the specific task of classification of genetic subtype, STRAP that employed Artistic Paintings and Natural Imaging as style sources outperformed STRAP that uses histopathologic imaging as style source. Despite promising performances, applying AdaIn as on-the-fly data augmentation is considered to be computationally expensive. Thus, stylized version of the datasets should be prepared in advance. A future perspective of this work is to apply this method to self-supervised learning, where it can be integrated as a data augmentation technique for contrastive learning framework.
@@
Vuong et al. [202] [^7] https://github.com/trinhvg/IMPash proposed a new augmentation strategy called PatchShuffling, which is inspired by Pretext-Invariant Representation Learning (PIRL) [133]. This strategy aims to enable DL model to learn invariant representation, resistant to domain shift. It is used during the pre-training phase, along with another type of augmentation InfoMin [192]. Unlike PIRL, which starts by extracting the patch feature and then rearranging these features within the initial image, PatchShuffling directly shuffles the initial image itself. Initially, PatchShuffling randomly selects a portion from the image, ensuring it size is approximately [0.6,0.1] of the original image area. This cropped image is then resized and randomly flipped. They randomly extract 9 non-overlapping patches and assemble them as 3-by-3 grid to form a new image. On the other hand, the InfoMin augmentation constructs two views of the original image: it is designed to minimize the mutual information between the original and the augmented version of an image, while preserving any task-relevant information intact. Their proposed framework showed promising results when tested for colorectal cancer tissue classification. It outperformed other traditional histology domain-adaptation and self-supervised learning methods. As a future step, the authors proposed to investigate the combination of PatchShuffling with other domain-specific augmentation.
@@
Xiong et al. [212] introduced Enhanced Domain Transformation (EDT) for improving DG on unseen images. The proposed technique is based on the idea that if a transformed color space represents the unseen data with a distribution identical to that of training data, as a result the model trained on that space should generalize more effectively. The EDT framework incorporates several image processing steps to enhance model generalization capabilities. Initially, it employs image local average subtraction, a technique inspired by Kaggle Diabetic Retinopathy (DR) winning solution [64]. Following this step, it applies average blurring for reducing high-frequency noise and Adaptive Local Contrast Enhancement (ALCE) for normalizing the images. Furthermore, it utilizes PCA color jittering which modifies the training image color with the predominant color component to simulate the color characteristics of the unseen domain. The provided image might originate from the known domain, unseen domain or even non-medical images (ImageNet, etc.). This method was applied for age regression and DR classification using fundus photographs. Despite promising results, the average blurring process can mask important features, reducing the model’s classification performance.

### v3

@@ 5.1.1 Data manipulation
Almahfouz Nasser et al. [29] and Yin et al. [30] proposed to use a pre-processing anto-encoder to reduce the domain shift problem. The main idea is to produce a uniform domain appearance of input images prior to applying a classification network. The auto-encoder is trained to reconstruct the input images using a Mean Square Error (MSE) loss. To further erase domain specific signals, adversarial learning is incorporated using a domain discriminator, which is a network tasked to detect the domain label (i.e., the scanner technology used to acquire images). On the other hand, the autoencoder is trained to maximize the domain label prediction loss and minimize the reconstruction loss simultaneously. Unfortunately, when applied for mitotic figure detection in Whole Slide Images (WSI) [29], this method performed very poorly with an F1 score of 0.0030 on the test set of the MIDOG 2021 challenge. Instead of applying the autoencoder in the spatial domain, Yin et al. [30] proposed to apply it in the frequency domain, under the assumption that the amplitude spectrum encodes the style information whereas the phase spectrum contains the content details. The goal was to learn a frequency attention map that can align different domain images in a common frequency domain. That is, the input image was first converted to the frequency domain. The phase spectrum of the input image remains unchanged. In contrast, its amplitude spectrum is reconstructed using an autoencoder which filters out domain specific frequency information. In the context of lung nodule detection from CT images, they reported a competition performance metric [104] of 0.911 on the target test set of LUNA-DG.
@@
Gunasinghe et al. [31] considered three classical preprocessing methods: median filter, input standardization, and randomized multi-image histogram matching. The median filter is a non-linear digital preprocessing technique, used to remove noise from an image. Input standardization, a method inspired by Quellec et al. [105], aims to attenuate illumination variations. Histogram matching is a technique that transforms the histograms of the red, green and blue channels of an image to match those of a specific reference image. In randomized multi-image histogram matching, histogram matching is performed sequentially using multiple reference images selected from the training source domain. When tested for glaucoma detection in fundus photographs using RIMONEv2 and REFUGE, the results have shown that standardization of images led to greater performances in most scenarios with an average Area Under the Receiver Operator Characteristic Curve (AUC) of 0.85.
Inspired by Nyúl et al. [106], Garrucho et al. [32] proposed to perform intensity scale standardization, a two-step technique consisting of: 1) a training step, where a standardized histogram is learned from the training images to identify key histogram landmarks, and 2) a transformation step, in which the images are adjusted using the parameters learned in the first step. When applied for mass detection in mammography, enhanced generalization performance were achieved, outperfoming MixStyle, Cutout, RandConv and histogram equalization.
@@
Wang et al. [33] proposed to use normalizing-flow-based method for counterfactual inference within a Structural Causal Model (SCM), to attain harmonization of data. The idea is to explicitly model the causal relationship of known confounders such as site, gender and age, and ROI features (i.e., the imaging measurement) in a SCM which uses normalizing flows to model probability distributions. Counterfactual inference can be performed upon such a model to sample harmonized data by intervening upon these variables. For the task of age regression and Alzheimer’s disease classification, this method showed better cross-domain generalization compared to state-of-the-art algorithms such as ComBat and IRM, and to models trained on raw data.
@@
In this context, Li et al. [19] proposed Amplitude Spectrum Diversification for single-DG to improve the diversity of training data. First, an input image is converted into the frequency domain using the Discrete Fourier transform. Then, diverse samples are generated by modifying the amplitude spectrum using a variety of randomization operations, i.e., randomize the amplitude and position of points in the amplitude spectrum using rescaling and pixel shuffling operations. One advantage of their proposed method is that no extra network is needed for adversarial sample generation. The authors reported an average accuracy over all out-of-domain data of 0.6285 for the MIDOG dataset and of 0.6287 on a multicenter colposcopic image dataset.
Zhang et al. [20] integrated a similar strategy, using domain randomization, which was implemented using amplitude mix or color jitter. In amplitude mix, an image is perturbed through linearly interpolating its amplitude spectrum with that of another image. In color jitter, variations are introduced in terms of hue, saturation and contrast distributions.
@@
Lucieri et al. [34] presented Amplitude-Focused Amplitude-Phase Recombination for pair samples (AF-APR-P) [^1] https://github.com/adriano-lucieri/shape-bias-in-dermoscopy. It aims to enhance the model’s ability to generalize by focusing on the amplitude spectrum of the images while altering the phase spectrum. This is achieved by swapping the phase spectrum among images but retaining their original amplitude spectrum. The authors showed improved performance on binary skin lesion classification tasks on the International Skin Imaging Collaboration (ISIC) dataset [^2] https://www.isic-archive.com/ and the seven-point checklist criteria dataset [107].
@@
Wang and Xia [35] extended the conventional mixup to cross-domain mixup to create a virtual domain based on the data from source domains. The original mixup technique produces convex combinations of pairs of images and their labels: it interpolates pairs of samples from the same domain that are drawn at random. In cross-domain mixup, one combine pairs of samples from different domains, to form a virtual domain $\mathcal{S}_{mix}$ that comprises virtual images ($x_{mix}$) and labels ($y_{mix}$), as formulated in the following equations:
@@
Experiments performed on chest X-rays datasets for the diagnosis of thoracic diseases showed that their proposed method outperformed Empirical Risk Minimization (ERM) and six other DG approaches.
@@
Garrucho et al. [32] also studied different augmentations techniques for DG. Namely, Cutout [108], RandConv [109] and MixStyle [110]. In addition, they investigated a data homogenization approach, the intensity scale standardization approach (presented in Section 5.1.1). They evaluated the performances of their model using one or a combination of data augmentation strategies. The experiments for mass detection in mammography showed that the combination of intensity scale standardization and cutout data augmentation led to the best results in all unseen domains.
@@
To enhance their model’s generalizability to different devices, Lafarge and Koelzer [36] incorporated a sequence of transformations such as transposition, color shift, Gamma correction, Hue rotation, spatial shift, additive Gaussian noise and cutout [108]. The evaluation of this method for mitotic figure detection on the preliminary test set of the MIDOG challenge resulted in a F1 score of 0.6828.
@@
Dexl et al. [37] performed a single augmentation to each image as part as a very simple random augmentation approach, inspired by Trivial Augment [111]. These augmentations are uniformly selected from a pool of color, noise, and special transformations, with the intensity of augmentation randomly chosen within a predefined range. Each image is also randomly flipped, and the RGB channels are shuffled at random. Using this strategy, their model achieved an F1 score of 0.7138 on the preliminary test phase of the MIDOG challenge.
@@
In the domain of histopathology, Long et al. [38] proposed to broaden the spectrum of stain color appearances in the training images. This was achieved by introducing randomness in selecting stain normalization techniques and target color styles. This method used two stain normalization techniques: Reinhard [112] and Vahadane [113]. Each technique was applied using a specific probability. To achieve robust detection performance for variety of images, they gradually expand the color style ranges to the network until there is a degradation in the detection performance. Their model achieved an F1 score of 0.7500 on the preliminary test phase of the MIDOG challenge.
@@
Li et al. [39], Chung et al. [40] and Scalbert et al. [41] proposed a GAN-based approach to expand the style variance of the training data. Li et al. [39] employed CycleGAN [^3] https://github.com/lizheren/MSVCL_MICCAI2021 to map the images of source domain to a device-style domain. The augmented images were then used in their contrastive learning strategies to develop a representation with better generalization capability to various device domains (Section 5.1.3). In addition, to enhance sample diversity, they used different diversifying operations including random cropping, random rotation, horizontal flipping, and adjustment of brightness, contrast, and saturation.
Chung et al. [40] adopted StarGAN[114] to translate images into arbitrary device styles (based on the mixing of device characteristics), without losing morphological information upon training. Next, a detection network was trained on the translated images for mitotic figure detection. Their model achieved an F1 score of 0.7548 on the preliminary test phase of the MIDOG challenge.
Scalbert et al. [41] introduced Test-Time data Augmentation (TTA) based on StarGANV2 [115] [^4] https://gitlab.com/vitadx/articles/test-time-i2i-translation-ensembling, a more recent multi-domain image-to-image translation model. The idea is to project images from unseen domain into each source domain, classify the generated images and ensemble their predictions. The proposed method has shown good results when evaluated for two different histopathology tasks: 1) patch classification of lymph node section WSIs and 2) tissue type classification in colorectal histological images. This method outperformed standard/ Hematoxylin&Eosin (H&E) specific color augmentation/normalization and standard test-time augmentation techniques.
@@
In their Style Transfer Augmentation for Histopathology (STRAP)  [^5] https://github.com/rikiyay/style-transfer-for-digital-pathology data augmentation approach, Yamashita et al. [42] proposed to use image-to-image translation models at the testing phase. They employed random style transfer from non-medical style source (such as natural images from the miniImageNet dataset [116]) by applying AdaIn style transfer [117], as in Geirhos et al. [118]. That is, the style of medical images (i.e., histopathology images), namely the texture, color and contrast are translated with the style of a selected non-medical image. However, the semantic content of the image, the global object shapes are unchanged. Their method was applied for 1) colorectal cancer classification into two distinct genetic sub-types based on WSI in a single-DG setting and 2) identifying the presence or absence of breast cancer metastases in image patches extracted from histopathlogic scans of lymph node sections in a multi-source DG setting. It achieved higher performances compared to stain normalization based approaches. Despite promising performances, applying AdaIn as on-the-fly data augmentation is considered to be computationally expensive.
@@
With the aim to learn invariant representation, resistant to domain shift, Vuong et al. [43] [^6] https://github.com/trinhvg/IMPash proposed a new augmentation strategy called PatchShuffling. Inspired by Pretext-Invariant Representation Learning (PIRL) [119], it is used during the pre-training phase, along with another type of augmentation InfoMin [120]. Unlike PIRL, which starts by extracting the patch feature and then rearranging these features within the initial image, PatchShuffling directly shuffles the initial image itself. Initially, PatchShuffling randomly selects a portion from the image, ensuring its size is approximately [0.6,0.1] of the original image area. This cropped image is then resized and randomly flipped. They randomly extract 9 non-overlapping patches and assemble them as 3-by-3 grid to form a new image. On the other hand, the InfoMin augmentation constructs two views of the original image: it is designed to minimize the mutual information between the original and the augmented version of an image, while preserving any task-relevant information intact. Their framework outperformed other traditional histology domain-adaptation and self-supervised learning methods in the task of colorectal cancer tissue classification.
@@
Xiong et al. [44] introduced Enhanced Domain Transformation (EDT) for improving DG on unseen images. It incorporates several image processing steps: 1) image local average subtraction, 2) average blurring for reducing high-frequency noise and adaptive local contrast enhancement for normalizing the images, 3) PCA color jittering which modifies the training image color with the predominant color component to simulate the color characteristics of the unseen domain. The provided image might originate from the known domain, unseen domain or even non-medical images (ImageNet, etc.). This method was applied for age regression and DR classification using fundus photographs. Despite promising results, the average blurring process can mask important features, reducing the model’s classification performance.